# 3D-Prover: Diversity Driven Theorem Proving With Determinantal Point Processes

### v1

@@ Abstract
A key challenge in automated formal reasoning is the intractable search space, which grows exponentially with the depth of the proof. This branching is caused by the large number of candidate proof tactics which can be applied to a given goal. Nonetheless, many of these tactics are semantically similar or lead to an execution error, wasting valuable resources in both cases. We address the problem of effectively pruning this search, using only synthetic data generated from previous proof attempts. We first demonstrate that it is possible to generate semantically aware tactic representations which capture the effect on the proving environment, likelihood of success, and execution time. We then propose a novel filtering mechanism which leverages these representations to select semantically diverse and high quality tactics, using Determinantal Point Processes. Our approach, 3D-Prover, is designed to be general, and to augment any underlying tactic generator. We demonstrate the effectiveness of 3D-Prover on the miniF2F-valid and miniF2F-test benchmarks by augmenting the ReProver LLM. We show that our approach leads to an increase in the overall proof rate, as well as a significant improvement in the tactic success rate, execution time and diversity.
@@ 1 Introduction
Interactive Theorem Proving, as the name suggests, has traditionally involved a human guiding a proving system to verify a formal proposition. It has found applications in a wide range of fields, from secure software (Tan et al., 2019) to the verification of mathematical results (Hales et al., 2017). There has been significant interest in automating this process, with formalization efforts requiring a high level of human expertise (Klein et al., 2009). Beyond this, it is considered a ‘grand challenge’ for AI, requiring a high level of reasoning and planning to be successful (Reddy, 1988). Even the largest current models struggle with the complexity of the task, with for example GPT-4 only able to solve 13.5% (Thakur et al., 2023) of the high school level miniF2F-test (Zheng et al., 2021) benchmark. This has motivated the development of specialized models and search algorithms to address the unique challenges of the domain (Wang et al., 2024; Polu et al., 2022; Jiang et al., 2022b; Han et al., 2022; First et al., 2023; Zhao et al., 2023; Wang et al., 2023; Whalen, 2016; Wu et al., 2021b; Wang et al., 2018; Wang & Deng, 2020; Rabe et al., 2020; Polu & Sutskever, 2020; Mikuła et al., 2023; Loos et al., 2017; Li et al., 2021; Lewkowycz et al., 2022; Jiang et al., 2021; 2022a; Gauthier et al., 2017).
With most non-trivial proofs requiring long chains of correct reasoning, it is a challenge to generate them in one pass without mistakes. The addition of a search algorithm is common for addressing this, as is done by the current state-of-the-art DeepSeek-Prover-V1.5 (Xin et al., 2024). Under this paradigm, candidate tactics are generated and executed in the proving system, which (if successful) results in new subgoals to prove. This generates a tree of possible proof paths, where a search algorithm selects the most promising nodes to expand. The primary challenge faced by these approaches is the exponential growth in the number of proof paths, limiting the complexity of the problems that can be solved efficiently.
Many of the generated tactics are equivalent, modulo variable renaming and other semantics-preserving transformations. See Figure 1 for a sample search tree from the ReProver (Yang et al., 2023) system, where several semantically similar paths are explored, wasting valuable resources. Simple lexical similarity scores fail to cover the semantics (meaning) of a tactic, as captured by the effect of the tactic on the environment. For example, an expression and its negation vary by only a single character, but have a large semantic difference. It is therefore desirable to filter tactics by their semantic rather than syntactic diversity. In addition, many tactics lead to an execution error from the prover. From our experiments with miniF2F, we find approximately 75% of tactics result in an execution error (Section 2.2). With the execution of tactics in the environment being expensive, this further restricts the space of proofs which can be explored efficiently.
@@
Figure 1: An example node expansion for a failed ReProver attempt, which our DPP model was able to prove. Tactics on the left result in the same proof state, tactics on the right result in an error, and tactics in the centre result in a unique proof state. The high error rate and tactic similarity motivates our filtering approach, which prunes the search space to give a diverse set of subgoals.
@@
These challenges motivate our proposed approach, Diversity Driven Determinantal Point Process Prover (3D-Prover). 3D-Prover adds an extra ‘dimension’ to existing proving systems by including a filtering mechanism on top of the existing tactic generation and search components. 3D-Prover uses Determinantal Point Processes (Kulesza, 2012) to prune the proof search space by filtering tactic candidates to a diverse and high quality subset. The large amount of synthetic data generated from proof attempts enables us to learn the effect tactics have on the environment, including the likelihood of an error and the execution time. We leverage this to generate tactic representations which reflect their semantics, which 3D-Prover uses to filter tactics based on a combination of their diversity and quality. 3D-Prover allows for a direct tradeoff between search objectives, with hyperparameters controlling the weighting of error, time and diversity in the filtering process. 3D-Prover is a general approach which can be used to augment any underlying tactic generator. We demonstrate this by augmenting the open source ReProver LLM to obtain a significant improvement in the success rate, execution time and diversity of tactics, and the overall proof success rate.
We summarize our contributions as follows:
@@
  • We propose a novel edge filtering approach using Determinantal Point Processes (Kulesza & Taskar, 2011), which leverage these representations to select semantically diverse subsets of quality tactics. Our method is modular and can be used with any underlying tactic model.
  • We evaluate our approach by augmenting ReProver (Yang et al., 2023) on the miniF2F (Zheng et al., 2021) benchmark, where we demonstrate a significant improvement in the tactic success rate, diversity and overall proof success rate.
@@ 1.1 Related work
There is little prior work on learning the effect of a tactic on the proving environment.  Xin et al. (2024) recently included successful environment responses as an auxiliary learning objective, however do not investigate the task in detail. We extend this by modeling the error likelihood, error messages and execution time, which we use to generate useful tactic representations. Several approaches have used previous proof attempts to improve performance, using the sparse binary signal from the proof status of a goal (Bansal et al., 2019; Wu et al., 2021a). This has been used to improve search algorithms, as done in (Lample et al., 2022; Polu et al., 2022; Wang et al., 2023). These approaches do not consider the diversity of the nodes expanded, with nothing preventing the search from exploring semantically similar paths.  Xin et al. (2024) uses intrinsic reward for exploration by rewarding new nodes in the search tree. The addition of any node is rewarded equally, even if they are similar (but not identical) to existing nodes. We instead select tactics based on their diversity with respect to the resulting nodes.  First & Brun (2022) use a diverse ensemble of models to improve proof performance, whereas we focus on diversity with respect to the environment response, given an arbitrary underlying model (or models).
1.2 Preliminaries
We first define the space of goals $\mathcal{S}$, tactics $\mathcal{T}$ and failures $\mathcal{F}$. For our purposes, these all contain arbitrary strings, with the goal being a formal proposition, the tactic a command and the failure an error message. A proof tree is a DAG $G=(V,E)$ where $V\subset\mathcal{S}$ is the set of goals and $E$ the edges between them. A proof attempt for a goal $g_{0}$ first initialises the proof tree with $V=\{g_{0}\},E=\emptyset$. The search policy $\pi_{S}:G\times V\rightarrow\mathbb{R}^{+}$ is a distribution over goals given a proof tree, being used to select a goal $g$ to expand. The tactic policy $\pi_{T}:\mathcal{S}\times\mathcal{T}\to\mathbb{R}^{+}$ is a distribution over tactics given a goal, where $N\in\mathbb{N}$ tactics are sampled to give tactics $\{t_{i}\}_{i=1}^{N}\subset\mathcal{T}$. The goal, tactic pairs $(g,t_{i})$ are then passed to the environment $\mathcal{E}:\mathcal{S}\times\mathcal{T}\to\mathcal{O}$. For each pair, after $\tau_{i}\in\mathbb{R}$ seconds, it returns either a new set of goals $g_{i}^{\prime}\subset\mathcal{S}$ or an error, $e_{i}\in\mathcal{F}$. We define this response as the output $o_{i}\in\mathcal{O}$ where ${\mathcal{O}}:=\mathcal{P}(\mathcal{S})\cup\mathcal{F}$. We further define the status $s_{i}\in\{0,1\}$ as 00 if $o_{i}\in\mathcal{F}$, $1$ if $o_{i}\in\mathcal{P}(\mathcal{S})$ and the transition as the tuple $(g,t_{i},s_{i},\tau_{i},o_{i})$. The proof tree is then updated with $G=G\cup g_{i}^{\prime}$ for all $g_{i}^{\prime}$, and the associated transitions are added as edges to $E$. This is repeated until a proof is found, or a budget is exhausted. A proof of $g$ is found when $\mathcal{E}(g,t_{i})=\emptyset$ for any $t_{i}$, or if all $\{g_{i}^{\prime}\}$ are proven for $\mathcal{E}(g,t_{i})=\{g_{i}^{\prime}\}\subset\mathcal{S}$. The result of a proof attempt is the set of transitions $\{(g_{k},t_{ki},s_{ki},\tau_{ki},o_{ki})\}$ for all selected goals $g_{k}$ and their expanded tactics $t_{i}$.
@@ 2 Transition Aware Representation Learning
Figure 2: Our architecture for learning transition aware tactic embeddings. The tokenized tactic $\bf{t}$ and goal $\bf{g}$ are concatenated and passed through the Encoder $E$. A single representation vector $\bf{e}$ is generated by mean-pooling over the tactic token embeddings $\bf{t^{\prime}}$. The Predictor $P$ takes this embedding and predicts whether the tactic results in an error (Status), and the execution time (Time). The Decoder $D$ takes the embedding and goal to predict the environment response (Output), which is either an error message or new goals to prove. The result is a compact representation of the tactic which captures its effect on the proving environment, enabling our proposed filtering model.
One proof attempt can generate a rather large amount of data. A single pass of the miniF2F-valid benchmark of 244 proofs results in approximately 500,000 transitions, capturing rich information about the error likelihood, execution time and resulting proof state or error message. This section explores the feasibility of using this transition data to learn how tactics affect the environment. We operationalise this as a supervised learning task: given a goal and tactic, we predict the error status, execution time and environment output. We effectively learn these targets from only this synthetic data, and further embed this information into a compact tactic representation. The upshot, as we show in Section 3, is that these representations can be utilised to improve the performance of subsequent proof attempts.
@@ 2.1 Transition Models
We assume a dataset $\mathcal{D}$ of transition tuples $\{(g_{k},t_{ki},s_{ki},\tau_{ki},o_{ki})\}$, as defined in 1.2. We learn a transition model $\xi:\mathcal{S}\times\mathcal{T}\to\{0,1\}\times\mathbb{R}\times\mathcal{O}$ which maps a goal $g_{k}$ and tactic $t_{ki}$ to an estimate of the status $s_{ki}$, time $\tau_{ki}$ and output $o_{ki}$. We construct our transition model $\xi$ with three components. For $d\in\mathbb{N}$, the Encoder $E:{\mathcal{S}}\times{\mathcal{T}}\to{\mathbb{R}}^{d}$ takes the goal $g_{k}$ and tactic $t_{ki}$ as input, and outputs a single embedding vector with unit norm, $E(g_{k},t_{ki})={\bm{e}}_{ki}$, $||{\bm{e}}_{ki}||=1$. The Predictor $P:{\mathbb{R}}^{d}\to[0,1]\times{\mathbb{R}}$ maps this embedding to a score for the time prediction and an error probability for the status, with $P({\bm{e}}_{ki})=(\hat{s}_{ki},\hat{\tau}_{ki})$. The Decoder $D:{\mathbb{R}}^{d}\times{\mathcal{S}}\to{\mathcal{O}}$ maps the embedding and goal to the output prediction, such that $D({\bm{e}}_{ki},g_{k})=\hat{o}_{ki}$. The transition model is then
$\xi(g_{k},t_{ki})=(P(E(g_{k},t_{ki})),D(E(g_{k},t_{ki}),g_{k}))=(P({\bm{e}}_{ki}),D({\bm{e}}_{ki},g_{k}))=(\hat{s}_{ki},\hat{\tau}_{ki},\hat{o}_{ki}).$ (1)
@@ 2.2 Experiments
For our experiments, we use an Encoder-Decoder Transformer for the Decoder $D$, and an Encoder-Only Transformer for the Encoder $E$. We take the pretrained ReProver (Yang et al., 2023) LLM to initialise both components. We implement the Predictor $P$ as a single hidden layer MLP, with hidden dimension $d/2$ (where $d=1472$) and two real valued output nodes. The time prediction $\hat{\tau}_{ki}$ is the output of the first node, and the status prediction $\hat{s}_{ki}$ is taken as the sigmoid of the second.
We investigate four instances of the transition model $\xi$. For the Combined model (Figure  2), the tactic is concatenated with the goal, and the embeddings from the Encoder are computed for all tokens. We then generate a single tactic embedding by mean-pooling over the tactic tokens. We compare this with the Separate model which encodes the tactic without attending to the goal. We hypothesise that allowing the tactic tokens to attend to the goal will allow the Encoder to better represent the semantics of the tactic. To form a naive baseline, we implement a No Tactic model which does not use the tactic at all, and instead uses only the goal tokens. We do this to account for any inherent patterns in the goal which may be predictive of the outcome, for example a particular goal which has a high error rate. This allows us to ground our results in the performance of this baseline, so we can observe the direct effect of the tactic in predictive performance. We also compare with an All Tokens model which uses all tactic tokens for the Decoder without reducing to a single embedding. We maintain the pooling operation over the tactic tokens for the status and time prediction tasks, but allow the Decoder to attend to all tokens for the output prediction. We implement this comparison to see the degree of information loss induced by reducing tactics to a single vector.
Given $\alpha_{s},\alpha_{\tau},\alpha_{o}\in\mathbb{R}^{+}$, with estimates $\hat{s}_{ki},\hat{\tau}_{ki},\hat{o}_{ki}$ and for a minibatch ${\mathcal{B}}\subseteq{\mathcal{D}}$, we optimise the transition loss
$\mathcal{L_{T}}({\mathcal{D}},\xi)=\sum_{(g_{k},t_{ki},s_{ki},\tau_{ki},o_{ki})\in{\mathcal{B}}}\alpha_{s}\mathcal{L}_{s}(s_{ki},\hat{s}_{ki})+\alpha_{\tau}\mathcal{L}_{\tau}(\tau_{ki},\hat{\tau}_{ki})+\alpha_{o}\mathcal{L}_{o}(o_{ki},\hat{o}_{ki}).$ (2)
@@
We obtain the dataset ${\mathcal{D}}$ from a vanilla ReProver attempt on miniF2F-valid, which results in 498,236 transitions, which we split randomly into 95% training, 5% testing. There is the possibility of dependence between the splits, as the test set includes goals seen in training with different tactics. The No Tactic baseline should capture any of this, with our results in Section 3.3.1 showing our representations generalise from miniF2F-valid to miniF2F-test. For the error prediction task, we reweight classes to account for imbalance, which is approximately 75% error, 25% success. We use the AdamW optimizer, with a learning rate of $10^{-5}$ and a batch size of 1. We train each model for 2 epochs on a single RTX4090, and report the results on the test set.
2.2.1 Results
@@
Table 1: Results for predicting unseen environment responses given a goal and tactic, for transitions from miniF2F-valid. The No Tactic result forms a baseline to assess the impact of the tactic representation. We observe that any tactic representation enables far better predictions, and constraining these to a single vector (Combined and Separate) does not hurt the performance gain. This demonstrates tactic representations which capture their effect on the environment, enabling our filtering model in Section 3. Comparing the Combined and Separate models, allowing the representation to attend to the goal leads to a large improvement.
To assess the Output prediction, we use beam search to generate 4 candidate outputs for each transition. We use the BLEU (Papineni et al., 2002) and ROUGE-L (Lin, 2004) scores to assess the quality of the highest scoring beam in comparison to the ground truth, which is either an error message or a new set of subgoals. We also report the Top-4 accuracy, which is the proportion of samples which have one beam identical to the ground truth. For the Status prediction task, we take the prediction as $1$ if $\hat{s}_{ki}>0.5$ and 00 otherwise, reporting the F1 score, true positive rate (TPR) and true negative rate (TNR). The Time MSE is the mean squared error of the time prediction over the test set.
@@
Table 1 summarises the performance of our transition models on the test set. Our results suggest tactic representations which capture useful information about their effect on the environment, which we can see by the clear improvement across all approaches compared to the No Tactic baseline. The higher scores across all metrics of the Combined versus the Separate model support our hypothesis that we can better predict transitions when the tactic embedding attends to the goal. The All Tokens model, where we allow the Decoder to attend to the full tactic, does not increase performance in comparison to the Combined model. This shows that we can effectively represent the tactic as a single embedding without any loss of information. Our results demonstrate the feasibility of learning the environment dynamics of proving systems, which is a difficult task. To illustrate this, we include all prediction examples for the Combined model, along with their ground truth, in the supplementary material.
3 Filtering Model
@@
Figure 3: DPP for tactic filtering. The tactic embeddings from the transition model are scaled by quality scores, before a subset of tactics are selected using $k$-DPP. Subsets are chosen proportionally to the area spanned by their elements, giving a combination of quality and diversity. For this simplified example, we take the 2D PCA projection of embeddings for tactics in Figure 1, setting the quality to the scaled generator logits. Comparing the shaded areas in (b) and assuming subst c and rw h1 have been selected, we see that symmetry is favoured over simp [h1]. Although simp [h1] is scored higher by the generator, it is less diverse with respect to subst c and rw h1.
In the previous section we used synthetic proof data to generate semantically aware tactic representations, allowing us to predict the likelihood of success, execution time and environment response. We now take this a step further by using these representations to augment proof search. We first present the necessary background on Determinantal Point Processes, enabling us to introduce our filtering model 3D-Prover, which prunes tactic candidates based on their quality and the semantic diversity of their representations. We show that 3D-Prover is able to improve the performance of the ReProver LLM on the miniF2F-valid and miniF2F-test benchmarks, particularly when a deeper search configuration is used. We conclude with a multifaceted ablation study showing the effect of our filtering model on the success rate, number of unique responses and execution time.
3.1 Determinantal Point Processes
Determinantal Point Processes (DPPs) are a class of probabilistic models for sampling subsets from a ground set $\mathcal{Y}$. In line with Kulesza (2012), for $|\mathcal{Y}|=n$ we define the kernel $L\in\mathbb{R}^{n\times n}$ of a DPP as the Gram matrix $L=B^{T}B$ for $B\in\mathbb{R}^{n\times d}$, where column ${\bm{b}}_{i}\in\mathbb{R}^{d}$ of $B$ is a vector representing element $i\in\{1,\dots,n\}$ of $\mathcal{Y}$. These vectors ${\bm{b}}_{i}$ are commonly decomposed into a set of unit norm diversity features $\bm{\phi}_{i}\in\mathbb{R}^{d}$ and quality scores $q_{i}\in\mathbb{R}^{+}$, so that ${\bm{b}}_{i}=q_{i}\bm{\phi}_{i}$, $||\bm{\phi}_{i}||=1$ for all $i\in\{1,\dots,n\}$. The similarity matrix $S$ is then defined as $S_{ij}=\bm{\phi}_{i}^{T}\bm{\phi}_{j}$.
The probability of sampling a subset $A\subseteq\mathcal{Y}$ from a DPP is then proportional to the determinant of the submatrix of $L$ indexed by $A$, $\mathbb{P}(A)\propto\text{det}(L_{A})=(\prod_{i\in A}q_{i}^{2})\text{det}(S_{A})$. Geometrically, this determinant is the volume of the parallelepiped spanned by the submatrix $L_{A}$, which as we see in Figure 3, is maximised based on a combination of the similarity and length (quality) of the chosen elements. In this way, DPPs elegantly trade off between the quality and diversity of elements. Normally the size of the sampled subset $|A|$ is variable, however Kulesza & Taskar (2011) introduce $k$-DPPs which restricts the size of the subset to a fixed $k\in\mathbb{N}$, and where the probability of sampling $A$ is normalised over subsets of size $k$. That is, for a $k$-DPP, $\mathbb{P}(A)\propto\text{det}(L_{A})/\sum_{|A^{\prime}|=k}\text{det}(L_{A^{\prime}})$.
3.2 Filtering Model
@@
Input :  Goal $g$, candidate tactics $T=\{t_{i}\}_{i=1}^{N}$, filter size $K$, Encoder $E$, Predictor $P$, error weight $\lambda_{s}$, time weight $\lambda_{\tau}$, temperature $\theta$, underlying tactic policy $\pi_{0}$

### v2

@@ Abstract
A key challenge in automated formal reasoning is the intractable search space, which grows exponentially with the depth of the proof. This branching is caused by the large number of candidate proof tactics which can be applied to a given goal. Nonetheless, many of these tactics are semantically similar or lead to an execution error, wasting valuable resources in both cases. We address the problem of effectively pruning this search, using only synthetic data generated from previous proof attempts. We first demonstrate that it is possible to generate semantically aware tactic representations which capture the effect on the proving environment, likelihood of success, and execution time. We then propose a novel filtering mechanism which leverages these representations to select semantically diverse and high quality tactics, using Determinantal Point Processes. Our approach, 3D-Prover, is designed to be general, and to augment any underlying tactic generator. We demonstrate the effectiveness of 3D-Prover on the miniF2F and LeanDojo benchmarks by augmenting popular open source proving LLMs. We show that our approach leads to an increase in the overall proof rate, as well as a significant improvement in the tactic success rate, execution time and diversity. We make our code available at https://github.com/sean-lamont/3D-Prover.
@@ 1 Introduction
Interactive Theorem Proving (ITP) traditionally involves a human guiding an ITP system to verify a formal proposition. The applications range from secure software (Tan et al., 2019) to the verification of mathematical results (Hales et al., 2017). There has been significant interest in automating this process, with formalization efforts requiring a high level of human expertise (Klein et al., 2009). It is also considered a ‘grand challenge’ for AI, requiring a high level of reasoning and planning to be successful (Reddy, 1988). Even large general purpose models struggle with the complexity of the task, with for example GPT-4 only able to solve 13.5% (Thakur et al., 2023) of the high school level miniF2F-test (Zheng et al., 2021) benchmark. This has motivated specialized models and search algorithms to address the unique challenges of the domain (see e.g.  (Li et al., 2024) for a review).
@@
Figure 1: An example node expansion for a failed ReProver attempt, which 3D-Prover was able to prove. Tactics on the left result in the same proof state, tactics on the right result in an error, and tactics in the centre result in a unique proof state. The high error rate and tactic similarity motivates our filtering approach, which prunes the search space to give a diverse set of subgoals.
With most non-trivial proofs requiring long chains of correct reasoning, it is a challenge to generate them in one pass without mistakes. The addition of a search algorithm is common for addressing this (Xin et al., 2024; Wu et al., 2024). Under this paradigm, candidate tactics are generated and executed in the proving system, which (if successful) results in new subgoals to prove. This generates a tree of possible proof paths, where a search algorithm selects the most promising nodes to expand. The primary challenge faced by these approaches is the exponential growth in the number of proof paths, limiting the complexity of the problems that can be solved efficiently. Many of the generated tactics are equivalent, modulo variable renaming and other semantics-preserving transformations. See Figure 1 for a sample search tree from ReProver (Yang et al., 2023), where several semantically similar paths are explored, wasting valuable resources. Simple lexical similarity scores fail to cover the semantics (meaning) of a tactic, as captured by the effect of the tactic on the environment. For example, an expression and its negation vary by only a single character, but have a large semantic difference. It is therefore desirable to filter tactics by their semantic rather than syntactic diversity. In addition, many tactics lead to an execution error from the prover. From our experiments with miniF2F, we find approximately 75-85% of tactics result in an execution error (Section 2.2). As tactic execution can be expensive, this further restricts the space of proofs which can be explored efficiently.
@@
These challenges motivate our proposed approach, Diversity Driven Determinantal Point Process Prover (3D-Prover). 3D-Prover adds an extra ‘dimension’ to existing proving systems by including a filtering mechanism on top of the existing tactic generation and search components. 3D-Prover uses Determinantal Point Processes (Kulesza, 2012) to prune the search space by filtering tactics to diverse and high quality subsets. The rich synthetic data generated from proof attempts enables us to learn the effect tactics have on the environment, including the error likelihood and execution time. We leverage this to generate tactic representations which reflect their semantics, which 3D-Prover uses to filter tactics based on a combination of their diversity and quality. 3D-Prover allows for a direct tradeoff between search objectives, with hyperparameters controlling the weighting of error, time and diversity in the filtering process. 3D-Prover is a general approach which can be used to augment any underlying tactic generator. We demonstrate this by augmenting the popular ReProver and InternLM2.5-Step-Prover LLMs to obtain a significant improvement in the success rate, execution time and diversity of tactics, and the overall proof success rate. To summarise our contributions:
@@
  • We propose a novel edge filtering approach using Determinantal Point Processes (Kulesza & Taskar, 2011), which leverage these representations to select semantically diverse subsets of quality tactics. Our method is modular and can be used with any tactic generator.
  • We evaluate our approach by augmenting ReProver (Yang et al., 2023) on the miniF2F (Zheng et al., 2021) benchmarks, where we demonstrate a significant improvement in the tactic success rate, diversity and overall proof rate.
@@ 1.1 Related work
There is little prior work on learning the effect of a tactic on the proving environment, with only  Xin et al. (2024) using successful environment responses as an auxiliary objective. We investigate the task in detail, as well as learning the error likelihood, error messages and execution time, which we use to generate useful tactic representations. Several approaches use previous proof attempts to improve performance, using the sparse binary signal from the proof result (Li et al., 2024). This has been used to improve search (Lample et al., 2022; Wang et al., 2023), however these approaches do not consider node diversity, with nothing preventing the exploration of semantically similar paths.  First & Brun (2022) examine a diverse ensemble of tactic models, whereas we focus on diversity with respect to the search, given an arbitrary underlying tactic model (or models). Recently, Yang et al. (2025) select subgoals based on their diversity, with a simple embedding model over the subgoal text. Our approach does not need to execute the tactics, as we learn embeddings reflecting the environment dynamics and use these to select tactics before execution, thereby saving resources.
1.2 Background: Determinantal Point Processes
Determinantal Point Processes (DPPs) are a class of probabilistic models for sampling subsets from a ground set $\mathcal{Y}$. They provide an inherent trade-off between the diversity and quality of the sampled subsets, successfully being applied to this end across a variety of domains (Kulesza, 2012; Hsiao & Grauman, 2018; Zhang et al., 2016). This motivates their use in our filtering approach (Section 4(a)).
In line with Kulesza (2012), for $|\mathcal{Y}|=n$ we define the kernel $L\in\mathbb{R}^{n\times n}$ of a DPP as the Gram matrix $L=B^{T}B$ for $B\in\mathbb{R}^{n\times d}$, where row ${\bm{b}}_{i}\in\mathbb{R}^{d}$ of $B$ represents element $i\in\{1,\dots,n\}$ of $\mathcal{Y}$. The ${\bm{b}}_{i}$ are commonly decomposed into a set of unit norm diversity features $\bm{\phi}_{i}\in\mathbb{R}^{d}$ and quality scores $q_{i}\in\mathbb{R}^{+}$, so that ${\bm{b}}_{i}=q_{i}\bm{\phi}_{i}$, $\bm{\phi}_{i}||=1$ for all $i\in\{1,\dots,n\}$. The similarity matrix $S$ is then defined as $S_{ij}=\bm{\phi}_{i}^{T}\bm{\phi}_{j}$. The probability of sampling $A\subseteq\mathcal{Y}$ is then proportional to the determinant of the submatrix of $L$ indexed by $A$, $\mathbb{P}(A)\propto\text{det}(L_{A})=(\prod_{i\in A}q_{i}^{2})\text{det}(S_{A})$. Geometrically, this determinant is the volume of the parallelepiped spanned by the submatrix $L_{A}$, which as we see in Figure 4(a), is maximised based on a combination of the similarity and length (quality) of the chosen elements. In this way, DPPs elegantly trade off between the quality and diversity of elements. Normally the size of the sampled subset $|A|$ is variable, however Kulesza & Taskar (2011) introduce $k$-DPPs which restricts the size of the subset to a fixed $k\in\mathbb{N}$, and where the probability of sampling $A$ is normalised over subsets of size $k$. That is, for a $k$-DPP, $\mathbb{P}(A)=\text{det}(L_{A})/\sum_{|A^{\prime}|=k}\text{det}(L_{A^{\prime}})$.
@@ 2 Transition Aware Representation Learning
One proof attempt can generate large amounts of data. We found a single pass of ReProver on the miniF2F-valid benchmark of 244 proofs results in approximately 500,000 transitions, capturing rich information about the error likelihood, execution time and resulting proof state or error message. We now explore the feasibility of using this data to learn how tactics affect the environment, operationalising this as a supervised learning task: given a goal and tactic, we predict the error status, execution time and environment output. We effectively learn these targets from only this synthetic data, and embed this information into a compact tactic representation. The upshot, as we show in Section 4(a), is that this can be used to improve the performance of subsequent proof attempts.
@@ 2.1 Transition Models
Figure 2: Our Combined architecture for learning transition aware tactic embeddings. The tactic $\bf{t}$ and goal $\bf{g}$ are concatenated and passed through the Encoder $E$. A representation vector $\bf{e}$ is generated by mean-pooling over the tactic token embeddings $\bf{t^{\prime}}$. The Predictor $P$ takes this embedding and predicts whether the tactic results in an error (Status), and the execution time (Time). The Decoder $D$ takes the embedding and goal to predict the environment response (Output), which is either an error message or new goals to prove. This setup yields a compact representation of the tactic which captures its effect on the proving environment, enabling our proposed filtering model.
The result of a proof attempt (formalised in A) is the dataset $\mathcal{D}$ of transitions $\{(g,t,s,\tau,o)\}$, which captures the results of applying tactics $t\in\mathcal{T}$ to goals $g\in\mathcal{S}$. The status $s\in\{0,1\}$, indicates a success (1) or failure (0), $\tau\in\mathbb{R}$ gives the execution time and the output $o\in\mathcal{O}$ is the environment response, which is an error message, new subgoals, or a proof success. We propose a method to learn tactic representations ${\bm{e}}\in\mathbb{R}^{d}$ which capture the result $(s,\tau,o)$ of applying $t$ to $g$. By using these as features for DPP, we can filter tactics based on their expected outcome, before they are executed.
[…]
We note that the Decoder and Predictor can only access information of $t$ through $\bm{e}$}. Hence our architecture requires the Encoder to learn an effective representation for $\bm{e}$}, so that the Decoder and Predictor can use this to determine the subsequent effect of the tactic on the environment.
@@ 2.2 Experiments
For our experiments, we use an Encoder-Decoder Transformer for the Decoder $D$, and an Encoder-Only Transformer for the Encoder $E$. We take the pretrained ReProver (Yang et al., 2023) LLM to initialise both components. We implement the Predictor $P$ as a single hidden layer MLP, with hidden dimension $d/2$ (where $d=1472$) and two real valued output nodes. The time prediction $\hat{\tau}$ is the output of the first node, and the status prediction $\hat{s}$ is taken as the sigmoid of the second. We use this simple Predictor architecture to speed up our filtering algorithm presented in Section 4(a).
We investigate several variations of the transition model $\xi$. For the Combined model (Figure  2), the tactic is concatenated with the goal, and the embeddings from the Encoder are computed for all tokens. We then generate a single tactic embedding by mean-pooling over the tactic tokens. We examine the Combined model both with the full goal text, and a variation Combined (Small Goal) which embeds the goal first, and concatenates it as a single token vector to the tactic. This variation allows for more efficient batching when used for filtering in Section 1, as the goal need only be embedded once for multiple tactics, however gives less information to the model. The Separate model encodes the tactic without attending to the goal. We hypothesise that allowing the tactic tokens to attend to the goal will allow the Encoder to better represent the semantics of the tactic. To form a naive baseline, we implement a No Tactic model which does not use the tactic at all, and instead uses only the goal tokens. We do this to account for any inherent patterns in the goal which may be predictive of the outcome, for example a particular goal which has a high error rate. This allows us to ground our results in the performance of this baseline, so we can observe the direct effect of the tactic in predictive performance. We also compare with an All Tokens model which uses all tactic tokens for the Decoder without reducing to a single embedding. We implement this comparison to see the degree of information loss induced by reducing tactics to a single vector. Given $\alpha_{s},\alpha_{\tau},\alpha_{o}\in\mathbb{R}^{+}$, with estimates $\hat{s},\hat{\tau},\hat{o}$ and for minibatch ${\mathcal{B}}\subseteq{\mathcal{D}}$, we optimise the following:
[…]
@@
We obtain $\mathcal{D}$} from a single ReProver attempt on miniF2F-valid, yielding 498,236 transitions split randomly into 95% training, 5% testing. For the error prediction task, we reweight classes to account for imbalance, which is approximately 75% error, 25% success. We use the AdamW optimizer, with a learning rate of $10^{-5}$ and a batch size of 1, train for 2 epochs, and report the results on the test set. To assess the Output prediction, we generate 4 outputs with beam search for each transition. We use BLEU (Papineni et al., 2002) and ROUGE-L (Lin, 2004) to assess the quality of the highest scoring beam in comparison to the ground truth, which is either an error message or new subgoals. The Top-4 accuracy is the proportion of samples with one beam identical to the ground truth. For Status, we take the prediction as $1$ if $\hat{s}_{ki}>0.5$ and 0 otherwise, reporting the F1 score, true positive rate (TPR) and true negative rate (TNR). For time, we take the Mean Squared Error (MSE) of the prediction.
Table 1: Results for predicting unseen environment responses given a goal and tactic, for transitions from miniF2F-valid. The No Tactic result forms a baseline to assess the impact of the tactic representation. We observe that any tactic representation enables far better predictions, and constraining these to a single vector (Combined and Separate) does not hurt the performance gain. This demonstrates tactic representations which capture their effect on the environment, enabling our filtering model in Section 4(a). Comparing the Combined and Separate models, allowing the representation to attend to the goal leads to a large improvement.
@@
Table 1 summarises the performance of our transition models. Our results suggest tactic representations which capture useful information about their effect on the environment  [^1] See Appendix E for further analysis of these embeddings., which we can see by the clear improvement across all approaches compared to the No Tactic baseline. The improvement of Combined over Separate supports our hypothesis that we can better predict transitions when the tactic embedding attends to the goal. As expected, Combined outperforms Combined (Small Goal), with Combined (Small Goal) significantly outperforming the Separate model. Combined (Small Goal) therefore gives an effective compromise between the accurate but expensive Combined model, and the goal-unaware Separate model. We note the All Tokens model, with the Decoder attending to the full tactic, does not improve upon the full Combined model. This shows our architecture effectively represents the tactic as a single embedding without losing any relevant information. These results are the first to demonstrate the feasibility of learning the environment dynamics of proof systems. To illustrate the difficulty of this task, all predictions for the Combined model and their ground truth are provided with our code.
3 3D-Prover
@@
(a) Visualisation of DPP for tactic filtering. The tactic embeddings from the transition model are scaled by quality scores, before a subset of tactics are selected using $k$-DPP. Subsets are chosen proportionally to the area spanned by their elements, giving a combination of quality and diversity. For this simplified example, we take the 2D PCA projection of embeddings for tactics in Figure 1, setting the quality to the scaled generator logits. Comparing the shaded areas in (b) and assuming subst c and rw h1 have been selected, we see that symmetry is favoured over simp [h1]. Although simp [h1] is scored higher by the generator, it is less diverse with respect to subst c and rw h1.
@@
Input :  Goal $g$, candidate tactics $T=\{t_{i}\}_{i=1}^{N}$, filter size $K$, Encoder $E$, Predictor $P$, error weight $\lambda_{s}$, time weight $\lambda_{\tau}$, temperature $\theta$, tactic policy $\pi_{0}$